OpenStudy (anonymous):

Find the domain (x^3+10x^2+31x+30)/x^2-4

OpenStudy (anonymous):

when will this function not exist? for what values of x will the function be undefined?

OpenStudy (anonymous):

\[\frac{ x^3+10x^2+31x+30 }{ x^2-4 }\]

OpenStudy (anonymous):

I have -2 and 2, but I don't understand why I don't include them.

OpenStudy (anonymous):

if for example you get X/0 or and number over zero you function will be undefined so you want to know when the denominator will equal zero

OpenStudy (anonymous):

Oh, duh, I remember that.

OpenStudy (anonymous):

if the denominator is zero your function is undefined so you set the denominator equal to zero x^2-4=0 x^2=4 x=+-2 so x=2 and x=-2 will both result in zero in the denominator which is undefined so these values arn't included in the domain

OpenStudy (anonymous):

So would I write that in interval notation or would this work: \[{x|x \neq-2,2}\]

OpenStudy (anonymous):

Also, if there's a case where I get an imaginary number, would I disregard it as a "problem" in the domain?

OpenStudy (anonymous):

When you have square roots or other even roots you need to know when the value under the radical is negative and exclude all of those values.

OpenStudy (anonymous):

complex values as solutions don't effect the domain.

OpenStudy (anonymous):

Thank you. You're a fab helper! :)